. Cookie Company produces parts for the auto industry. Part 2 is machined in Department #1, which

has the following budgeted conversion costs. Use 2,400 production hours and .03 hours per unit to

manufacture.

Labor

$220,000

Depreciation

30,000

Maintenance

10,000

Supplies

25,600

Total

$285,600

(a) Compute Department #1’s budgeted cell conversion cost rate for the current year.

(b) Compute Part 2’s budgeted cell conversion cost per unit.
Answer #1

a) Budgeted cell conversion Cost rate = 285600/2400 = 119 per hour

b) Budgeted cell conversion Cost per unit = 285600/80000 = 3.57 per unit